1.)A ball starts at rest and rolls down a ramp, developing a speed of 50 m/s in 5 seconds. Calculate the ball's acceleration.
2.)A car slows to a stop with an average acceleration of - 3.0 m/s2. How long will it take to slow the car from 15 m/s to a position of rest?

Answer by Alan3354(69443)

1.)A ball starts at rest and rolls down a ramp, developing a speed of 50 m/s in 5 seconds. Calculate the ball's acceleration.
a = v/t
a = 50 m/s / 5 sec
a = 10 m/sec/sec
=============================
2.)A car slows to a stop with an average acceleration of - 3.0 m/s2. How long will it take to slow the car from 15 m/s to a position of rest?
v = at
t = v/a = 15/3
t = 5 seconds
-------------
Ignore the minus sign, it only indicates the speed is decreasing as time increases.